Lilian9667:

Solve: 5x-7x+6=-2(x-3 0 3/4 2 infinitely many soulutiions

Lilian9667:

May I get help :)

mxddi3:

ok combine like terms on the left. 5x-7x= ?

darkknight:

Also distribute on the right.

darkknight:

-2(x-3) is...?

mxddi3:

^ i was gonna get there but she never answered lol

Lilian9667:

sorry

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight -2(x-3) is...? \(\color{#0cbb34}{\text{End of Quote}}\) Is that a Equation I'm confused

darkknight:

mxddi3 u wanna take it?

umm:

@Lilian9667, what they're saying is, does your equation look like this? \[5x-7x+6=-2(x-3)\]

mxddi3:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 \(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight -2(x-3) is...? \(\color{#0cbb34}{\text{End of Quote}}\) Is that a Equation I'm confused \(\color{#0cbb34}{\text{End of Quote}}\) you have to distibute, so what is -2 times x? and what is -2 x -3?

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @umm @Lilian9667, what they're saying is, does your equation look like this? \[5x-7x+6=-2(x-3)\] \(\color{#0cbb34}{\text{End of Quote}}\) yes

Lilian9667:

-2x + 6 = -2x + 6 0x = 0

umm:

First thing's first. We'll start out by simplify both sides of this equation.\[5x-7x+6=-2(x-3)\] Distribute. \[5x+-7x+6=(-2)(x)+(-2)(-3)\] \[5x+-7x+6=-2x+6\] Combine like terms. \[(5x+-7x)+(6)=-2x+6\] \[-2x+6=-2x+6\]\[-2x+6=-2x+6\] Add 2x to both sides. \[-2x+6+2x=-2x+6+2x\] \[6=6\] Subtract 6 from both sides. \[6-6=6-6\] Which comes out to your final answer. \[0=0\]

darkknight:

yeah lilian, u did that right

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight yeah lilian, u did that right \(\color{#0cbb34}{\text{End of Quote}}\) Thanks, And THANK ya'll for helping me =D

umm:

Anytime. Good luck.

darkknight:

np